OpenStudy (anonymous):

Which number equals the square root of 1, 792.? 14 4the square root of 7. 16the square root of 7. 7the square root of 14.

OpenStudy (misty1212):

HI!! how about a nice calculator?

OpenStudy (anonymous):

i dont have one...

OpenStudy (misty1212):

lol of course you do dear, you are on a computer you have an almost unlimited number of calculators i prefer this one http://www.wolframalpha.com/input/?i=sqrt%281792%29

OpenStudy (misty1212):

ooh look the answer is C what a shock C, is the preferred answer by over 5 to 1

jimthompson5910 (jim_thompson5910):

Why the answer is choice C \[\Large \sqrt{1792} = \sqrt{256*7}\] \[\Large \sqrt{1792} = \sqrt{256}*\sqrt{7}\] \[\Large \sqrt{1792} = 16\sqrt{7}\] notice how I factored 1792 into two numbers where one of the numbers are a perfect square (the largest perfect square factor possible) Then I used the rule \(\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\)

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

oh thanks(:

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]